# Incentivize Contribution and Learn Parameters Too: Federated Learning with Strategic Data Owners

### v2

@@ Abstract
Classical federated learning (FL) assumes that the clients have a limited amount of noisy data with which they voluntarily participate and contribute towards learning a global, more accurate model in a principled manner. The learning happens in a distributed fashion without sharing the data with the center. However, these methods do not consider the incentive of an agent for participating and contributing to the process, given that data collection and running a distributed algorithm is costly for the clients. The question of rationality of contribution has been asked recently in the literature and some results exist that consider this problem. This paper addresses the question of simultaneous parameter learning and incentivizing contribution, which distinguishes it from the extant literature. Our first mechanism incentivizes each client to contribute to the FL process at a Nash equilibrium and simultaneously learn the model parameters. However, this equilibrium outcome can be away from the optimal, where clients contribute with their full data and the algorithm learns the optimal parameters. We propose a second mechanism with monetary transfers that is budget balanced and enables the full data contribution along with optimal parameter learning. Large scale experiments with real (federated) datasets (CIFAR-10, FeMNIST, and Twitter) show that these algorithms converge quite fast in practice, yield good welfare guarantees, and better model performance for all agents.
@@ 1 Introduction
A high quality machine learning model is built when the model is trained on a large amount of data. However, in various practical situations, e.g., for languages, images, disease modeling, such data is split across multiple entities. Moreover, in many applications, data lies in edge devices and a model is learnt when these edge devices interact with a parameter server. Federated Learning (FL) [18, 23, 13, 39] is a recently developed distributed learning paradigm where the edge devices are users’ personal devices (like mobile phones and laptops) and FL aims to leverage on-device intelligence. In FL, there is a center (also known as the parameter server) and several agents (edge devices). To learn the model parameters, the edge devices can only communicate through the center. However, a major challenge of federated learning is data-scarcity, i.e., owing to its limited storage capacity, the edge devices possesses only a few amounts of data, which may not be sufficient for the learning task. Hence, the end user participates in a distributed learning process, where she exploits the data of similar users present in the system.
In the classical federated learning problem, it is typically assumed that the agents participate with all of their data-points voluntarily. However, in the presence of rational users, such an assumption may not always hold since sampling data is costly for agents. It may happen that some agents try to contribute very few data-points and try to exploit the system by learning based on others’ data-points. This phenomenon is called free-riding [14] and disincentivizes honest participants to contribute their data to the FL process. Hence, naturally, a few recent works have concentrated on designing incentive mechanisms in federated learning [25, 6, 2].
@@
“Can a mechanism be designed that incentivizes each user to participate and contribute their maximum data in FL and simultaneously learn the optimal model parameters?”
In this paper, we address this question in two stages. First, we consider a utility model of the agents where the accuracy of learning a parameter and the contribution levels of all the agents give every agent some benefit, while their individual contribution levels lead to their personal costs. We propose an algorithm, namely Updated Parameter Best Response Dynamics (UPBReD) in the federated learning setting that achieves simultaneous contribution and learning of the model parameters by the agents. However, this method suffers from a sub-optimal contribution by the agents and therefore the global learning is also sub-optimal. In the second stage, we improve this dynamics by allowing monetary transfer where contributors and consumers are treated differently using the monetary transfer and incentivizes all agents to contribute their maximum amount of data to the learning process. This mechanism, namely Two Phase Updated Parameter Best Response Dynamics (2P-UPBReD), learns the optimal parameters for all agents and asks the data-consumers to pay and pays the data-contributors, without keeping any surplus. 2P-UPBReD is distinguished in the fact that it simultaneously learns the optimal model parameters, incentivizes full contribution from the agents, yet is quite simple to use in practice.
1.1 Related work
Federated Learning [18] has gained significant attention in the last decade or so. The success story of FL is primarily attributed to the celebrated FedAvg algorithm [23], where one of the major challenges of FL, namely communication cost is reduced by local steps. In subsequent works, several other challenges of FL, such as data heterogeneity [16, 10], byzantine robustness [37, 15, 12], communication overhead [29, 17, 11] and privacy [34, 33] were addressed.
Later works focused on incentivizing client participation with advanced aggregation methods. Approaches included dynamic weighting of client updates to ensure that the global model outperforms local models [5], Shapley value to assess contributions [32], contract theory to maximize fairness in data contribution [14], and fairness by eliminating malicious clients and rewarding those who enhance performance of the model [8].
Monetary incentives in FL have been explored through various mechanisms. Yu et al. [38] proposed dynamic budget allocation to maximize utility and reduce inequality. Blockchain-based methods [27] incentivize high-quality data contributions under budget constraints, and client rewards have been shown to improve final model utility [36]. Georgoulaki and Kollias [9] analyzed utility-sharing in FL games, showing a price of anarchy of two and price of stability of one under budget-balanced payments. Auction-based FL systems use reinforcement learning to let agents adjust bids for profit while preserving accuracy [30], and centers can optimize budgets to enhance utility and reduce delays [31].
Coalitional approaches have also been studied. Donahue and Kleinberg [6] modeled FL as a coalitional game allowing joint training, while Ray Chaudhury et al. [28] introduced CoreFed to ensure core stability. Their work was generalized to ordinal utility settings [4], and Shapley value-based approaches ensure reciprocal fairness [24]. However, these do not consider non-cooperative agents choosing data contributions strategically. Repeated-game formulations [21] show that subgame perfect equilibria can be inefficient, prompting budget-balanced mechanisms that ensure social efficiency and individual rationality. Incentive-aware learning frameworks by Blum et al. [2] and their extension with best response dynamics and budget-balance by Murhekar et al. [25] achieve Nash equilibria and maximum welfare under constraints. Yet, these works overlook simultaneous model training, a gap our work addresses by incorporating strategic data contribution and model parameter learning together.
@@ 1.2 Our contributions
  • We propose a mechanism called Updated Parameter Best Response Dynamics, UPBReD (Algorithm 1) that allows simultaneous learning and contribution by the agents in FL. This mechanism does not use any monetary transfers.
@@
  • We show that UPBReD converges to a pure strategy Nash equilibrium (Theorem 1).
@@
  • However, the Nash equilibrium can be different from the socially optimal outcome where agents contribute with their full dataset $s^{\max}$ and the center learns the optimal model parameters $w^{\texttt{{OPT}}}$ (Example 1).
@@
  • We then propose an updated and cleaner two-phase mechanism namely Two Phase Updated Parameter Best Response Dynamics, 2P-UPBReD (Algorithm 2) that allows monetary transfers only among the agents, i.e., budget balanced. The data contributors (those who contribute above average quantities of data) get paid and the data consumers (those who contribute below average quantities of data) make the payments. However, both types of agents learn the optimal model parameters.
@@
  • We show that the Nash equilibrium for 2P-UPBReD now shifts to where all agents make their maximum contribution of data $s^{\max}$ and learn the optimal model parameters $w^{\texttt{{OPT}}}$ (Theorem 2).
@@
  • Experiments on real datasets (CIFAR-10, FeMNIST, and Twitter) demonstrate that while UPBReD leads to suboptimal social welfare, 2P-UPBReD achieves performance comparable to FedAvg for all datasets (see Section 5). Notably, FedAvg performs poorly in strategic settings.
@@ 2 Preliminaries
Consider a federated learning setup where a set of data contributors, given by $N=\{1,2,\ldots,n\}$, is interacting with a center. Each data contributor (agent) $i$ has access to a private labeled dataset $D_{i}$, with the size of the dataset given by $s_{i}^{\max}=|D_{i}|$. The agents are interested in learning a parameter vector $w\in\mathbb{R}^{m}$ from these data so that it helps them predict some unlabeled data accurately (e.g., to perform a classification task). However, each individual agent has a limited amount of data, and learning $w$ only from that data may not be accurate enough. So, they learn this parameter via a federated learner such that the model is trained on the consolidated data of all the $n$ users. Assume that the datasets are drawn from the same distribution, e.g., all the agents sample human disease data from a certain geographical location. However, sampling such data is costly and agent $i$ incurs a cost $c_{i}(s_{i})$ when she trains the model locally on her dataset of size $s_{i}\in S_{i}:=[0,s_{i}^{\max}]$. Under this setup, each agent $i$ gets a utility based on how much data $s_{i}$ she decides to sample and train on. Hence, the utility is given by
$u_{i}(w,s_{i},s_{-i})=a_{i}(w,s_{i},s_{-i})-z_{i}(s_{i},s_{-i}),$ (1)
@@
where $s_{-i}$ is the data chosen by the agents other than $i$ in this federated learning process. The function $a_{i}(w,s_{i},s_{-i})$ is the accuracy function, which denotes the benefit to agent $i$ if the parameter learned by the center is $w$ and the agents contribute by running the federated learning algorithm on their dataset sizes given by the vector $s=(s_{i},s_{-i})$.
{remark}
@@
[Arguments of the $u_{i}$ function] Note that we have explicitly assumed that the utility function $u_{i}$ depends on $(w,s_{i},s_{-i})$. We now motivate this dependence. In FL problems, we typically run iterative algorithms to optimize the accuracy (or loss) function. Note that, with this, the weight $w^{t}$ at time $t$ depends on all the $s_{i}^{\ell}$, $\ell<t,\forall i\in N$. However, the utility $u_{i}$ at time $t$ depends on $(s_{i}^{t},s_{-i}^{t})$ as well, which is not captured through $w^{t}$. Hence, we require the said explicit dependencies.
The function $z_{i}(s_{i},s_{-i})$ is the effective cost to agent $i$ which may depend on the data contributions of all agents. In the first part of this paper, we consider the case where the cost is only personal to agent $i$, i.e., $z_{i}$ is only a function of agent $i$’s chosen data size $s_{i}$. In the second part, we assume that this cost may be additionally taxed or subsidized by the center based on their contributed data-sizes.
@@
Since the agents are strategic, every agent $i$’s aim is to maximize her utility by appropriately choosing her strategy $s_{i}$ given the strategies $s_{-i}$ of the other players and the parameter $w$ chosen by the center. The center, on the other hand, is interested in learning the parameter $w$ that maximizes the sum of the accuracies as follows, when all agents contribute their maximum data-sizes, i.e., $s_{i}^{\max},i\in N$.
$w^{\texttt{{OPT}}}\in\operatorname*{\arg\!\max}_{w}\sum_{i\in N}a_{i}(w,s_{i}^{\max},s_{-i}^{\max}).$ (2)
@@
Note that the goal of the center does not consider the effective costs of the agents since those are incurred by the agents. We will refer to the term $\sum_{i\in N}a_{i}(w,s_{i},s_{-i})$ as the social welfare in this context. We are interested in the question that whether we can design a federated learning algorithm that can make $(s_{i}^{\max},s_{-i}^{\max})$ a Nash equilibrium of the underlying game and the center can learn $w^{\texttt{{OPT}}}$. In this context, the Nash equilibrium is defined as follows.
@@
This definition is a modification of the standard definition of Nash equilibrium since the equilibrium profile $(s_{i}^{*},s_{-i}^{*})$ depends on the parameter $w$ (we do not write it explicitly for notational cleanliness). Existence of such a PSNE in this context is obvious due to Nash’s theorem [26]. To see this, consider another game $\Gamma^{\prime}$ where the pure strategies are to pick 00 or $s_{i}^{\max}$ for every $i\in N$. Every mixed strategy of $\Gamma^{\prime}$ is a pure strategy of the original data contribution game we consider above. Since by Nash’s theorem, a mixed strategy Nash equilibrium exists for any finite game, PSNE exists in our game. We will refer to the term PSNE as Nash equilibrium (NE) in the rest of the paper.
We also aim for the following property of a mechanism that involve monetary transfers.
@@
This property ensures that the net monetary in or out-flow is zero and the mechanism only allows monetary redistribution among the agents. In the following section, we consider the general case where the effective cost is arbitrary.

### v3

@@ Abstract
Classical federated learning (FL) assumes that the clients have a limited amount of noisy data with which they voluntarily participate and contribute towards learning a global, more accurate model in a principled manner. The learning happens in a distributed fashion without sharing the data with the center. However, these methods do not consider the incentive of an agent for participating and contributing to the process, given that data collection and running a distributed algorithm is costly for the clients. The question of rationality of contribution has been asked recently in the literature and some results exist that consider this problem. This paper addresses the question of simultaneous parameter learning and incentivizing contribution in a truthful manner, which distinguishes it from the extant literature. Our first mechanism incentivizes each client to contribute to the FL process at a Nash equilibrium and simultaneously learn the model parameters. We also ensure that agents are incentivized to truthfully reveal information in the intermediate stages of the algorithm. However, this equilibrium outcome can be away from the optimal, where clients contribute with their full data and the algorithm learns the optimal parameters. We propose a second mechanism that enables the full data contribution along with optimal parameter learning. Large scale experiments with real (federated) datasets (CIFAR-10, FEMNIST, and Twitter) show that these algorithms converge quite fast in practice, yield good welfare guarantees and better model performance for all agents.
@@ 1 Introduction
A high quality machine learning model is built when the model is trained on a large amount of data. However, in various practical situations, e.g., for languages, images, disease modeling, such data is split across multiple entities. Moreover, in many applications, data lies in edge devices and a model is learnt when these edge devices interact with a parameter server. Federated Learning (FL) (Konečnỳ, 2016; McMahan et al., 2017; Kairouz et al., 2021; Zhang et al., 2021) is a recently developed distributed learning paradigm where the edge devices are users’ personal devices (like mobile phones and laptops) and FL aims to leverage on-device intelligence. In FL, there is a center (also known as the parameter server) and several agents (edge devices). To learn the model parameters, the edge devices can only communicate through the center. However, a major challenge of federated learning is data-scarcity, i.e., owing to its limited storage capacity, each edge device possesses only a limited amount of data, which may not be sufficient for the learning task. Hence, the end user participates in a distributed learning process, where they exploits the data of similar users present in the system. As a result of their participation the users gain some benefit from the shared model that is learned and incur some loss for their contributions to the learning process.
In the classical federated learning problem, it is typically assumed that the agents participate with all of their data-points voluntarily. However, in the presence of rational users, such an assumption may not always hold since sampling data is costly for agents. It may happen that some agents try to contribute very few data-points and try to exploit the system by learning based on others’ data-points. This phenomenon is called free-riding (Karimireddy et al., 2022) and disincentivizes honest participants to contribute their data to the FL process. Hence, naturally, a few recent works have concentrated on designing incentive mechanisms in federated learning (Murhekar et al., 2023; Donahue and Kleinberg, 2021; Blum et al., 2021).
@@
“Can a mechanism be designed that incentivises each user to participate and contribute their maximum data in FL and simultaneously learn the optimal model parameters?”
In this paper, we address this question in two stages. First, we consider a utility model of the agents where the quality of the learned parameter and the contribution levels of all the agents give every agent some benefit, while their individual contribution levels lead to their personal costs. We propose an algorithm, namely Updated Parameter Best Response Dynamics (UPBReD) in the federated learning setting that achieves simultaneous contribution and learning of the model parameters by the agents. We introduce payments to ensure that agents truthfully share their contributions. However, this method suffers from a sub-optimal contribution by the agents and therefore the global learning is also sub-optimal. In the second stage, we improve this dynamics with monetary transfers where contributors and consumers are treated differently using the monetary transfer and incentivizes all agents to contribute their maximum amount of data to the learning process. This mechanism, namely Two Phase Updated Parameter Best Response Dynamics (2P-UPBReD), learns the optimal parameters for all agents and asks the data-consumers to pay and pays the data-contributors, without keeping any surplus. 2P-UPBReD is distinguished in the fact that it simultaneously learns the optimal model parameters, incentivizes full contribution from the agents, yet is quite simple to use in practice.
@@ 1.1 Our contributions
  • We propose a mechanism called Updated Parameter Best Response Dynamics, UPBReD (Algorithm˜1) that allows simultaneous learning and contribution by the agents in FL. This mechanism does not use any monetary transfers.
@@
  • We show that UPBReD converges to a pure strategy Nash equilibrium (Theorem˜1).
  • We add payments to UPBReD to construct a mechanism T-UPBReD that elicits truthful reports from agents in every round (Theorem˜2).
@@
  • However, the Nash equilibrium can be different from the socially optimal outcome where agents contribute with their full dataset $s^{\max}$ and the center learns the optimal model parameters $w^{\texttt{{OPT}}}$ (˜3).
@@
  • We then propose an updated and cleaner two-phase mechanism namely Two Phase Updated Parameter Best Response Dynamics, 2P-UPBReD (Algorithm˜3) that allows monetary transfers only among the agents, i.e., budget balanced. The data contributors (those who contribute above average quantities of data) get paid and the data consumers (those who contribute below average quantities of data) make the payments. However, both types of agents learn the optimal model parameters.
@@
  • We show that the Nash equilibrium for 2P-UPBReD now shifts to where all agents make their maximum contribution of data $s^{\max}$ and learn the optimal model parameters $w^{\texttt{{OPT}}}$ (Theorem˜3).
@@
  • Experiments on real datasets (CIFAR-10, FEMNIST, and Twitter) demonstrate that while UPBReD leads to suboptimal social welfare, 2P-UPBReD achieves performance comparable to FedAvg for all datasets (see Section˜5). Notably, FedAvg performs poorly in strategic settings.
1.2 Related work
Federated learning (Konečnỳ, 2016) has gained significant attention in the last decade or so. The success story of FL is primarily attributed to the celebrated FedAvg algorithm (McMahan et al., 2017), where one of the major challenges of FL, namely communication cost is reduced by local steps. In subsequent works, several other challenges of FL, such as data heterogeneity (Karimireddy et al., 2020b; Ghosh et al., 2020a), byzantine robustness (Yin et al., 2018; Karimireddy et al., 2020a; Ghosh et al., 2021), communication overhead (Stich et al., 2018; Karimireddy et al., 2019; Ghosh et al., 2020b) and privacy (Wei et al., 2020; Truex et al., 2020; Kumar et al., 2025) were addressed.
Later works focused on incentivizing client participation with advanced aggregation methods. Approaches included dynamic weighting of client updates to ensure that the global model outperforms local models (Cho et al., 2022), Shapley value to assess contributions (Tastan et al., 2024), contract theory to maximize fairness in data contribution (Karimireddy et al., 2022), and fairness by eliminating malicious clients and rewarding those who enhance performance of the model (Gao et al., 2021),  (Procaccia et al., 2025) aim to maximize utilitarian welfare in the heterogenous setting under PAC learning constraints.
Monetary incentives in FL have been explored through various mechanisms. Yu et al. (2020) proposed dynamic budget allocation to maximize utility and reduce inequality. Blockchain-based methods (Pandey et al., 2022) incentivize high-quality data contributions under budget constraints, and client rewards have been shown to improve final model utility (Yang et al., 2023)Liu et al. (2020) use a blockchain based system to enable Shapley value based profit distributions. Georgoulaki and Kollias (2023) analyzed utility-sharing in FL games, showing a price of anarchy of two and price of stability of one under budget-balanced payments. Auction-based FL systems use reinforcement learning to let agents adjust bids for profit while preserving accuracy (Tang and Yu, 2023), and centers can optimize budgets to enhance utility and reduce delays (Tang and Yu, 2024).
Coalitional approaches have also been studied. Donahue and Kleinberg (2021) modeled FL as a coalitional game allowing joint training, while Ray Chaudhury et al. (2022) introduced CoreFed to ensure core stability. Their work was generalized to ordinal utility settings (Chaudhury et al., 2024), and Shapley value-based approaches ensure reciprocal fairness (Murhekar et al., 2024). However, these do not consider non-cooperative agents choosing data contributions strategically. Repeated-game formulations (Mao et al., 2024) show that subgame perfect equilibria can be inefficient, prompting budget-balanced mechanisms that ensure social efficiency and individual rationality. Truthful cost reports are explored by Bornstein et al. (2024). Chakarov et al. (2025) introduce a budget balanced payment rule that is Bayesian incentive compatible in the heterogenous setting. Incentive-aware learning frameworks by Blum et al. (2021) and their extension with best response dynamics and budget-balance by Murhekar et al. (2023) achieve Nash equilibria and maximum welfare under constraints. Yet, these works overlook simultaneous model training, a gap our work addresses by incorporating strategic data contribution and model parameter learning together.
@@ 2 Preliminaries
Consider a federated learning setup where a set of data contributors, given by $N=\{1,2,\ldots,n\}$, is interacting with a center. Each data contributor (agent) $i$ has access to a private labeled dataset $D_{i}$, with the size of the dataset given by $s_{i}^{\max}=|D_{i}|$. The agents are interested in learning a parameter vector $w\in\mathbb{R}^{m}$ from these data so that it helps them predict some unlabeled data accurately (e.g., to perform a classification task). However, each individual agent has a limited amount of data, and learning $w$ only from that data may not be accurate enough. So, they learn this parameter via a federated learner such that the model is trained on the consolidated data of all the $n$ users. Assume that the datasets are drawn from the same distribution, e.g., all the agents sample human disease data from a certain geographical location. However, sampling such data is costly and agent $i$ incurs a cost, given by a function $c_{i}(s_{i})$, when it trains the model locally on its dataset of size $s_{i}\in S_{i}:=[0,s_{i}^{\max}]$. Along with these costs there may be some other arbitrary costs that an agent incurs based on the mechanism and other agent contributions. Under this setup, each agent $i$ gets a utility based on how much data $s$ has been chosen by the agents to sample and train on. Hence, the utility is given by
$u_{i}(w,s_{i},s_{-i})=v_{i}(w,s_{i},s_{-i})-z_{i}(s_{i},s_{-i}),$ (1)
@@
where $s_{-i}$ is the data chosen by the agents other than $i$ in this federated learning process. The function $v_{i}(w,s_{i},s_{-i})$ is the valuation function, which denotes the benefit to agent $i$ if the parameter learned by the center is $w$ and the agents contribute by running the federated learning algorithm on their dataset sizes given by the vector $s=(s_{i},s_{-i})$. The function $z_{i}(s_{i},s_{-i})$ is the effective cost to agent $i$ which may depend on the data contributions of all agents.
Federated learning is an iterative process, the center updates a parameter $w^{t}$ at round $t$, agents use this to choose their dataset size $s_{i}^{t+1}$, compute and share gradients $d_{i}^{t+1}\in\mathbb{R}^{m}$ with the center. The center aggregates gradients to update the parameter to $w^{t+1}$. We use the superscript $t$ when referring to these terms in a specific iteration of the training round.
@@
Remark (Arguments of the $u_{i}$ function). Note that we have explicitly assumed that the utility function $u_{i}$ depends on the parameter learned and the data contributions $(w,s_{i},s_{-i})$. We now motivate this dependence. In FL problems, we typically run iterative algorithms to optimize the accuracy (or loss) function. Note that, with this, the weight $w^{t}$ at time $t$ depends on all the $s_{i}^{\ell}$, $\ell<t,\forall i\in N$. However, the utility $u_{i}$ at time $t$ depends on $(s_{i}^{t},s_{-i}^{t})$ as well, which is not captured through $w^{t}$. Hence, we require the said explicit dependencies. We represent this dependency via two functions: the valuation function $v_{i}$ gives the benefit the agent gets from the parameter learned and the data contributions, while the effective cost is only dependent on all players’ data contributions.
Remark (Realizing the $v_{i}$ function). $v_{i}(w,s)$ denotes the benefit agent $i$ derives from the learned parameters at a given contribution vector. It can be realistically realised by an agent using the negative of loss computed on their data for the given parameter $w$. For experiments in Section˜5 we use the cross entropy loss.
@@
Since the agents are strategic, every agent $i$’s aim is to maximize its utility by appropriately choosing its strategy $s_{i}$ given the strategies $s_{-i}$ of the other players and the parameter $w$ chosen by the center. The center, on the other hand, is interested in learning the optimal parameter $w$ that maximizes the sum of the valuations as follows, when all agents contribute their maximum data-sizes, i.e., $s_{i}^{\max},i\in N$.
$w^{\texttt{{OPT}}}\in\operatorname*{\arg\!\max}_{w}\sum_{i\in N}v_{i}(w,s^{\max})$ (2)
@@
Note that the goal of the center does not consider the effective costs of the agents since those are incurred by the agents. We will refer to the term $\sum_{i\in N}v_{i}(w,s_{i},s_{-i})$ as the social welfare in this context. We are interested in the question of whether we can design a federated learning algorithm that can make $(s_{i}^{\max},s_{-i}^{\max})$ a Nash equilibrium of the underlying game and the center can learn $w^{\texttt{{OPT}}}$. In this context, the Nash equilibrium is defined as follows.
@@
This definition is a modification of the standard definition of Nash equilibrium since the equilibrium profile $(s_{i}^{*},s_{-i}^{*})$ depends on the parameter $w$ (we do not write it explicitly for notational cleanliness). Existence of such a PSNE in this context is obvious due to Nash’s theorem (Nash Jr, 1950). To see this, consider another game $\top$ where the pure strategies are to pick 0 or $s_{i}^{\max}$ for every $i\in N$. Every mixed strategy of $\top^{\prime}$ is a pure strategy of the original data contribution game we consider above. Since by Nash’s theorem, a mixed strategy Nash equilibrium exists for any finite game, PSNE exists in our game. We will refer to the term PSNE as Nash equilibrium (NE) in the rest of the paper.
We also aim for the following property of a mechanism that involves monetary transfers.
@@
This property ensures that the net monetary in or out-flow is zero and the mechanism only allows monetary redistribution among the agents.
[…]
Definition 3 (Ex-Post Incentive Compatible (EPIC)).
A mechanism with a given allocation rule $a$ and a payment rule $p$ is EPIC at round $t$ if for every agent $i\in N$ for the true type and valuation profile $\theta,v$ and a misreported type and valuation profile $\hat{\theta},\hat{v}$,
$u_{i}(\hat{\theta}_{i}^{t},\theta_{-i}^{t},\hat{v}_{i}^{t},v_{-i}^{t}|\theta^{t},v^{t})\leqslant u_{i}({\theta}^{t},{v}^{t}|\theta^{t},v^{t}).$
EPIC ensures that an agent gains no improvement in utility for misreporting when other agents are reporting truthfully. This is the strongest truthfulness notion for mechanism design with interdependent valuations (valuation of an agent depends on the types of all agents) since satisfying DSIC is impossible in such settings (see (Mezzetti, 2004) for further details). In the following section, we consider the general case where the effective cost is arbitrary.